OpenStudy (anonymous):

if cos theta =-4/7 what are the values of sin theta and tan theta

OpenStudy (anonymous):

i did it halfway through but stuck at the fraction part

OpenStudy (ddcamp):

What method are you trying to use?

OpenStudy (anonymous):

sin 2 theta + cos 2 theta = 1

OpenStudy (ddcamp):

So, you have: \[\sin^2(\theta) + (\frac{4}{7})^2=1?\] What is \[(\frac{4}{7})^2\]

OpenStudy (anonymous):

yes and then i got sin 2 theta + 16/49

OpenStudy (anonymous):

I subtract on both sides but i dont know how to subtract fraction from whole

OpenStudy (ddcamp):

Here's a clue: \[1=\frac{49}{49}\]

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

1/49?

OpenStudy (anonymous):

ive never seen that b4 lol

OpenStudy (ddcamp):

\[\frac{49}{49} - \frac{16}{49} = \frac{49-16}{49}\]

OpenStudy (ddcamp):

So, \[\sin^2(\theta) = \frac{33}{49} \rightarrow \sin(\theta) = \frac{\sqrt{33}}{\sqrt{49}} = \frac{\sqrt{33}}{7}\]

OpenStudy (anonymous):

Ohh okay I see! Ok Ill tell u wat my final answer is

OpenStudy (anonymous):

sin theta will be sine theta =+/- sqrt 33/7

OpenStudy (anonymous):

can u help with one more. Its very confusin

OpenStudy (anonymous):

@ddcamp